# Do ImageNet-trained models learn shortcuts? The impact of frequency shortcuts on generalization

### v1

@@ Abstract
Frequency shortcuts refer to specific frequency patterns that models heavily rely on for correct classification. Previous studies have shown that models trained on small image datasets often exploit such shortcuts, potentially impairing their generalization performance. However, existing methods for identifying frequency shortcuts require expensive computations and become impractical for analyzing models trained on large datasets. In this work, we propose the first approach to more efficiently analyze frequency shortcuts at a larger scale. We show that both CNN and transformer models learn frequency shortcuts on ImageNet. We also expose that frequency shortcut solutions can yield good performance on out-of-distribution (OOD) test sets which largely retain texture information. However, these shortcuts, mostly aligned with texture patterns, hinder model generalization on rendition-based OOD test sets. These observations suggest that current OOD evaluations often overlook the impact of frequency shortcuts on model generalization. Future benchmarks could thus benefit from explicitly assessing and accounting for these shortcuts to build models that generalize across a broader range of OOD scenarios. Codes are available at link.
@@ 1 Introduction
Superficial correlations between data and ground truth [9, 5] can be learned by models to minimize the training objectives with the least effort [36]. This learning behavior is called shortcut learning, which either harms the generalization performance of models or gives an illusion of good generalization abilities when learned shortcuts are present in OOD test sets [41]. Artifacts and visual cues that cause shortcut learning are recognizable by visual inspection of images, and their impact on the training of models can be mitigated through data selection or augmentation [29, 5, 21], i.e. counteracting the spurious correlations between data and ground truth.
However, there exist shortcuts in the Fourier domain, which are implicitly embedded in image data characteristics and not easily detectable by visual inspection [41, 39]. Such shortcut solutions consist of small frequency subsets that are easy-to learn and sufficient for models to achieve high classification rate. Typically, they correspond to simple features like textures, shapes or colors in the spatial domain. Wang, et al. [41] identified frequency shortcuts by retaining relevant frequencies to classification of a certain class. The relevance of an individual frequency was measured by the loss value of the model tested on images of the class with that frequency removed. Thus, their approach requires computational time increasing proportionally to the number of classes in a dataset and the image resolution. Yet, frequency shortcuts usually consist of multiple frequencies. The method in [41] might overlook some shortcuts as individual frequency relevance to classification neglects joint contribution of frequencies. Furthermore, the computational burden limits its applicability for analyzing shortcut learning behavior of models trained on larger datasets (e.g. ImageNet) with hundreds or thousands of classes.
@@
ImageNet-1k 8500 (354 days) 198 (8.25 days)
@@
    We discover that ImageNet-trained models (both CNN and transformer architectures) are subject to learn frequency shortcuts. Different from shortcuts formed by visual cues [9], frequency shortcuts (easy-to-learn features) lead to good performance on both in-distribution (ID) and OOD tests if they do not block the models from learning other semantics (difficult features).
@@ 2 Related works
Models can learn shortcut solutions based on superficial correlations between data and ground truth [9, 11] to optimize training objectives with the least effort. This learning behavior is due to simplicity-bias [36], which is caused by inductive biases provided by gradient descent or components like ReLUs [37]. Shortcuts can be visual cues in the data like source tags, artificial markers, etc. [21, 29]. For instance, one fifth of the horse images from Pascal VOC dataset were found to contain a source tag, on which models rely as discriminant feature to recognize horses [21]. Next to these visual cues, vision models are also subject to shortcuts implicitly existing in the frequency domain, which manifest as small sets of frequencies contributing significantly to image classification performance [41, 39]. The study in [41] explains why models exhibit a textures-bias for classification [8] from a frequency shortcut perspective, but does not provide further investigation on how frequency shortcuts affect the generalization and robustness performance of ImageNet-trained models, due to computational burden.
@@
Uncovering shortcuts helps understanding the generalizability of models on different out-of-distribution (OOD) data, explaining why models fail to generalize to or perform well on OOD data. It can provide shortcut-related prior knowledge to develop techniques that improve model generalization and robustness performance based on shortcut mitigation [40, 29]. However, common approaches are limited to identifying visually inspectable shortcuts such as text, watermark and color patches [21, 29, 26], using e.g. saliency maps [35]. Rather than directly identifying shortcuts in the data, in [5, 2] shortcut features present in individual images were quantified by assessing how difficult they were for models to learn. The authors in [30, 32, 34] uncovered shortcut features learned within representation space. The authors in [25] investigated shortcut learning by analyzing the relationship between semantic concepts using a knowledge graph, sharing a similar idea to [45]. Despite visual shortcut cues, there are non-observable shortcuts in the frequency domain, which are embedded in data characteristics. The work of [41] identifies relevant frequencies to classification, which potentially contain shortcut information. As this technique measures the relevance of one frequency at a time, it requires computational costs increasing proportionally to the number of classes and image resolution. Thus, limited attention has been given to uncovering frequency shortcuts in large datasets, with existing analyses primarily focused on datasets containing a small number of classes [41, 2, 45, 39].
Existing methods to identify shortcuts are limited to either manual inspection [21], or involve a time-consuming algorithm [41, 39]. Methods quantifying shortcut information or uncovering shortcut features primarily aim at mitigating shortcut learning, without explaining how shortcuts in the data impact model generalization performance. We are the first to enable frequency shortcut analysis of models trained on large-scale datasets and link them to generalization performance in different OOD settings.
@@ 3 Method
We propose a method to identify (non-visible) frequency shortcuts linked to intrinsic data characteristics rather than visual cues as in [29]. We reduce significantly computational demands compared to [41]. Our method is the first solution for shortcut identification that enables the analysis models trained on datasets with a large number of samples and classes. In the following sections, we detail our measurement of shortcut learning in models and examine the impact of shortcuts on generalization. Codes will be public.
@@ 3.1 Identifying frequency shortcuts
In order to discover the frequency subsets that models rely heavily on for correct classification of each class, we sample different frequency combinations that potentially contains frequency shortcuts. We use random sampling as it allows to easily consider joint contribution of frequencies to classification, improving the effectiveness of frequency shortcut identification. Random sampling has been shown to contribute to stable search results in data augmentation and reinforcement learning [23, 28, 31], while requiring lower computations w.r.t. optimization-based methods.
@@
In summary, the overall process of HFSS is as follows. Each stage of HFSS consists of three steps, namely (1) sampling $B$ frequency subsets which contain $p\%$ frequency patches, (2) evaluating shortcut information, and (3) forwarding subsets for next stage. At the initial search stage, we sample frequency patches with a large size and evaluate their shortcut information. We evaluate the frequency subsets according to their contributions to classification: we process a subset of training images by retaining a frequency subset and measure the class-wise loss of a model. Frequency subsets that contribute to a lower loss value are stronger candidates to indicate frequency shortcuts as they are sufficient for the model to achieve a high prediction score. We rank frequency subsets by the loss value and use the top-$N$ masks of each class in the next search stage.
Starting from the second search stage, we sample frequency patches one of the top-$N$ frequency subsets in the previous stage, with a smaller patch size. This confines the search space. In the final search stage, the top-$1$ frequency subset of each class is considered dominating the classification of the class concerned. Same as [41], we use binary masks to represent these frequency subsets, called Dominant Frequency Maps ($\mathrm{DFMs}$).
@@ 3.2 Measuring the degree of shortcut learning
Analyzing models learning behavior in a class-wise manner, as done in [41], is labor-intensive if there are hundreds or thousands of classes. To provide a broad overview of model learning behavior on training datasets, we categorize all classes in a dataset into two groups: (1) classes subject to shortcuts, and (2) non-shortcut classes. A class is considered subject to shortcuts if its true positive rate (TPR) surpasses a given threshold when the model is tested on images filtered to retain only dominant frequencies of the class concerned. We use TPR, as a high TPR indicates that the frequency subset is sufficient to achieve a high classification rate, thereby acting as a shortcut. The frequency-filtered images are referred to as DFM-filtered images.
For each class, we test the model on both the original test images and DFM-filtered images, computing TPR for both. We denote the TPR on the original images of class $c_{i}$ as $\mathrm{TPR}_{c_{i}}$ and the TPR on the DFM-filtered images as $\mathrm{TPR}^{DFM}_{c_{i}}$. If $\mathrm{TPR}^{DFM}_{c_{i}}{>}t$ (where $t{\in}[0,1]$ is a predefined threshold), class $c_{i}$ is considered subject to shortcuts; otherwise, it is considered as one of the non-shortcut classes. We compute the average $\mathrm{TPR}$ values of both groups, shortcut and non-shortcut classes, at different thresholds $t$. We note them as $\mathrm{AvgTPR_{sct}@t}$, $\mathrm{AvgTPR^{DFM}_{sct}@t}$ for the shortcut classes, and $\mathrm{AvgTPR_{non-sct}@t}$ and $\mathrm{AvgTPR^{DFM}_{non-sct}@t}$ for the non-shortcut classes. A higher threshold $t$ corresponds to analyzing stronger shortcuts, as dominant frequencies alone are sufficient for accurate classification. An example of the results of these metrics at different threshold values are given in Fig. 3, where the size of each point reflects the number of classes in the two groups. Larger points indicate a higher number of classes, which suggests that the model has a stronger tendency to shortcut learning.
@@
Figure 3: ResNet18 tested on original test images and DFM-filtered images of ImageNet-1k. Blue line shows the average TPR of classes subject to shortcuts on the original test images, and the orange line shows the results of non-shortcut classes, at different threshold $t$. The green and red lines correspond to results tested on DFM-filtered images. The number of shortcut classes depends on $t$, where lower $t$ indicates weak shortcuts and higher $t$ signifies stronger ones. The size of each point reflects the number of classes. The larger the size, the more classes included.
@@ 3.3 Impact of shortcut on generalization
To assess how frequency shortcuts impact the generalization and robustness performance of models trained on ImageNet-1k, we evaluate their performance on several datasets. We perform ID tests on ImageNet-1k (IN-1k) [6] and ImageNet-v2 (IN-v2) [33] test sets. They are considered in-distribution as they are collected using the same protocol as the training set. Furthermore, we use ImageNet-C (IN-C) [15] to benchmark the robustness of models against appearance image corruptions, e.g. noise, blur and weather changes. To evaluate the generalizability of models to images with different renditions, we use ImageNet-Renditions (IN-R) [16] and ImageNet-Sketch (IN-S) [38]. IN-R contains renditions like cartoons, paintings, art, toys of 200 classes from IN-1k. IN-S has the same number of classes as IN-1k, containing images of hand-drawn sketches. These datasets contain visual renditions, which serve to evaluate how the reliance on texture cues (corresponding to most frequency shortcuts) impacts on generalization. To measure the adversarial robustness of models, we apply the fast gradient sign method (FGSM) attacks [10] to the validation set of IN-1k, with $L_{\infty}=4/255$. We calculate the average TPR of shortcut and non-shortcut classes on OOD data, that we compare to the results on ID data.
@@ 4.1 Configuration of HFSS
Figure 7: Average TPR of shortcut and non-shortcut classes given different thresholds on datasets with 1000 classes. Models generally perform better on images of shortcut classes than non-shortcut classes. This does not hold for results on IN-S, which lacks preserved texture information. Circles are annotated with the number of shortcut classes.
Figure 8: Average TPR of shortcut and non-shortcut classes at different thresholds on datasets with 200 classes. Models perform better on shortcut classes than non-shortcut classes in IN-200. IN-R lacks preserved texture information and thus the models have worse performance on shortcut classes than non-shortcut classes. Circles are annotated with the number of shortcut classes.
@@ 4.2 Shortcut identification on IN-1k
Different from the results on IN-v2 and IN-C, frequency shortcuts impair generalization of models to texture and rendition changes. Model performance on shortcut and non-shortcut classes on IN-S is close to each other. In Fig. 8, we observe that shortcut classes have worse prediction results than non-shortcut classes when tested on IN-R which contains images with rendition changes. This is attributable to the fact that shortcut information are not available in the OOD tests, as the rendition and sketch test sets preserve less or very different texture information than IN-C and IN-v2 (see green lines of IN-R and IN-200, IN-S and IN-1k in Fig. 4 and Fig. 8). Frequency shortcut learning is an explicit cause of the texture-bias of ImageNet models, and the resulting impaired generalizability to renditions is in line with [8].
@@
Figure 9: An image of dugong contains frequency shortcuts which a model relies on for high-confidence classification. Its presence in OOD test sets like IN-C, and in the image under FGSM attacks yield correct predictions. But the reliance on shortcuts impairs model generalization performance to renditions changes.
@@ 4.3 Comparison with the existing approach
Frequency shortcuts can impair the generalization performance of models or provide a false impression of good generalization when shortcuts exist in OOD tests [41]. We thus validate the impact of shortcuts searched by HFSS in generalization tests, by measuring performance results on IN-SCT, an OOD test set designed by [41]. We report the results in Tab. 3. The strong shortcuts for classes airliner, frog and container ship searched by HFSS are present in the OOD set, contributing to close-to or above average TPR of classes Military aircraft (Mil-aircraft), tree frog, and fishing vessel in the IN-SCT dataset. The shortcuts searched by [41] for classes Siamese cat result in a much lower TPR for class tabby cat in IN-SCT, though shortcuts for this class still exist in the OOD data (as observed from the results of HFSS). This indicates that the method proposed in [41] has limitations in estimating the impact of shortcuts on OOD performance, and further demonstrates the effectiveness of HFSS in finding shortcuts.
@@
Table 3: $\mathrm{TPR}$ results on IN-SCT (first row) and corresponding DFM-filtered images (second and third rows). $\mathrm{TPR}$ above or close to average TPR (0.367) is highlighted in bold.

### v2

@@ Abstract
Frequency shortcuts refer to specific frequency patterns that models heavily rely on for correct classification. Previous studies have shown that models trained on small image datasets often exploit such shortcuts, potentially impairing their generalization performance. However, existing methods for identifying frequency shortcuts require expensive computations and become impractical for analyzing models trained on large datasets. In this work, we propose the first approach to more efficiently analyze frequency shortcuts at a large scale. We show that both CNN and transformer models learn frequency shortcuts on ImageNet. We also expose that frequency shortcut solutions can yield good performance on out-of-distribution (OOD) test sets which largely retain texture information. However, these shortcuts, mostly aligned with texture patterns, hinder model generalization on rendition-based OOD test sets. These observations suggest that current OOD evaluations often overlook the impact of frequency shortcuts on model generalization. Future benchmarks could thus benefit from explicitly assessing and accounting for these shortcuts to build models that generalize across a broader range of OOD scenarios. Codes are available at https://github.com/nis-research/hfss.
@@ 1 Introduction
Superficial correlations between data and ground truth [9, 5] can be learned by models to minimize training objectives with the least effort [36]. This learning behavior is called shortcut learning, which either harms the generalization performance of models or gives an illusion of good generalization abilities when learned shortcuts are present in OOD test sets [41]. Artifacts and visual cues that cause shortcut learning are recognizable by visual inspection of images, and their impact on the training of models can be mitigated through data selection or augmentation [29, 5, 21], i.e. counteracting the spurious correlations between data and ground truth.
However, there exist shortcuts in the Fourier domain, which are implicitly embedded in image data characteristics and not easily detectable by visual inspection [41, 39]. Such shortcut solutions consist of small frequency subsets that are easy-to-learn and sufficient for models to achieve high classification rate. Typically, they correspond to simple features like textures, shapes or colors in the spatial domain. Wang, et al. [41] identified frequency shortcuts by retaining relevant frequencies to classification of a certain class. The relevance of an individual frequency was measured by the loss value of the model tested on images of the class with that frequency removed. Thus, their approach requires computational time increasing proportionally to the number of classes in a dataset and the image resolution. Yet, frequency shortcuts usually consist of multiple frequencies. The method in [41] might overlook some shortcuts as individual frequency relevance to classification neglects joint contribution of frequencies. Furthermore, the computational burden limits its applicability for analyzing shortcut learning behavior of models trained on large datasets (e.g. ImageNet) with hundreds or thousands of classes.
@@
ImageNet-1k 8500 (354 days) 220 (9.2 days)
@@
    We discover that ImageNet-trained models (both CNN and transformer architectures) are subject to learn frequency shortcuts. Different from shortcuts formed by visual cues [9], frequency shortcuts (easy-to-learn features) lead to good performance on both in-distribution (ID) and OOD tests if they do not block models from learning other semantics (difficult features).
@@ 2 Related works
Models can learn shortcut solutions based on superficial correlations between data and ground truth [9, 11] to optimize training objectives with the least effort. This learning behavior is due to simplicity-bias [36], which is caused by inductive biases provided by gradient descent or components like ReLUs [37]. Shortcuts can be visual cues in the data like source tags and artificial markers [21, 29]. For instance, one fifth of the horse images from Pascal VOC dataset were found to contain a source tag, on which models rely as a discriminant feature to recognize horses [21]. Next to these visual cues, vision models are also subject to shortcuts implicitly existing in the frequency domain, which manifest as small sets of frequencies contributing significantly to image classification performance [41, 39]. The study in [41] explains why models exhibit a textures-bias for classification [8] from a frequency shortcut perspective, but does not provide further investigation on how frequency shortcuts affect the generalization and robustness performance of ImageNet-trained models, due to computational burden.
@@
Uncovering shortcuts helps understanding the generalizability of models on different out-of-distribution (OOD) data, explaining why models fail to generalize to or perform well on OOD data. It can provide shortcut-related prior knowledge to develop techniques that improve model generalization and robustness performance based on shortcut mitigation [40, 29]. However, common approaches are limited to identifying visually inspectable shortcuts such as text, watermark and color patches [21, 29, 26], using e.g. saliency maps [35]. Rather than directly identifying shortcuts in the data, in [5, 2] shortcut features present in individual images were quantified by assessing how difficult they were for models to learn. The authors in [30, 32, 34] uncovered shortcut features learned within representation space. The authors in [25] investigated shortcut learning by analyzing the relationship between semantic concepts using a knowledge graph, sharing a similar idea to [45]. Despite visual shortcut cues, there are non-observable shortcuts in the frequency domain, which are embedded in data characteristics. The work of [41] identified relevant frequencies to classification, which potentially contain shortcut information. As this technique measures the relevance of one frequency at a time, it requires computational costs increasing proportionally to the number of classes and image resolution. Thus, limited attention has been given to uncovering frequency shortcuts in large datasets, with existing analyses primarily focused on datasets containing a small number of classes [41, 2, 45, 39].
Existing methods to identify shortcuts are limited to either manual inspection [21] or involve a time-consuming algorithm [41, 39]. Methods quantifying shortcut information or uncovering shortcut features primarily aim at mitigating shortcut learning, without explaining how shortcuts in the data impact model generalization performance. We are the first to enable frequency shortcut analysis of models trained on large-scale datasets and link them to generalization performance in different OOD settings.
@@ 3 Method
We propose a method to identify (non-visible) frequency shortcuts linked to intrinsic data characteristics rather than visual cues as in [29]. We reduce significantly computational demands compared to [41]. Our method is the first solution for shortcut identification that enables the analysis of models trained on datasets with a large number of samples and classes. In the following sections, we detail our measurement of shortcut learning in models and examine the impact of shortcuts on generalization.
@@ 3.1 Identifying frequency shortcuts
In order to discover the frequency subsets that models rely heavily on for correct classification of each class, we sample different frequency combinations that potentially contain frequency shortcuts. We use random sampling as it allows to easily consider joint contribution of frequencies to classification, improving the effectiveness of frequency shortcut identification. Random sampling has been shown to contribute to stable search results in data augmentation and reinforcement learning [23, 28, 31], while requiring lower computations w.r.t. optimization-based methods.
@@
In summary, the overall process of HFSS is as follows. Each stage of HFSS consists of three steps, namely (1) sampling $B$ frequency subsets which contain $p\%$ frequency patches, (2) evaluating shortcut information, and (3) forwarding subsets for next stage. At the initial search stage, we sample frequency patches with a large size and evaluate their shortcut information. We evaluate the frequency subsets according to their contributions to classification: we process a subset of training images by retaining a frequency subset and measure the class-wise loss of a model. Frequency subsets that contribute to a lower loss value are stronger candidates to indicate frequency shortcuts as they are sufficient for the model to achieve a high prediction score. We rank frequency subsets by the loss values and use the top-$N$ masks of each class in the next search stage.
Starting from the second search stage, we sample frequency patches from one of the top-$N$ frequency subsets in the previous stage, with a smaller patch size. This confines the search space. In the final search stage, the top-$1$ frequency subset of each class is considered dominating the classification of the class concerned. Same as [41], we use binary masks to represent these frequency subsets, called Dominant Frequency Maps ($\mathrm{DFMs}$).
@@ 3.2 Measuring the degree of shortcut learning
Analyzing models learning behavior in a class-wise manner, as done in [41], is labor-intensive if there are hundreds or thousands of classes. To provide a broad overview of model learning behavior on training data, we categorize all classes in a dataset into two groups: (1) classes subject to shortcuts and (2) non-shortcut classes. A class is considered subject to shortcuts if its true positive rate (TPR) surpasses a given threshold when the model is tested on images filtered to retain only dominant frequencies of the class concerned. We use TPR, as a high TPR indicates that the frequency subset is sufficient to achieve a high classification rate, thereby acting as a shortcut. The frequency-filtered images are referred to as DFM-filtered images.
For each class, we test the model on both the original test images and DFM-filtered images, computing TPR for both. We denote the TPR on the original images of class $c_{i}$ as $\mathrm{TPR}_{c_{i}}$ and the TPR on the DFM-filtered images as $\mathrm{TPR}^{DFM}_{c_{i}}$. If $\mathrm{TPR}^{DFM}_{c_{i}}{>}t$ (where $t{\in}[0,1]$ is a predefined threshold), class $c_{i}$ is considered subject to shortcuts; otherwise, it is considered as one of the non-shortcut classes. We compute the average $\mathrm{TPR}$ values of both groups, shortcut and non-shortcut classes, at different thresholds $t$. We note them as $\mathrm{AvgTPR_{sct}@t}$, $\mathrm{AvgTPR^{DFM}_{sct}@t}$ for the shortcut classes, and $\mathrm{AvgTPR_{non-sct}@t}$ and $\mathrm{AvgTPR^{DFM}_{non-sct}@t}$ for the non-shortcut classes. A higher threshold $t$ corresponds to stronger shortcuts, as dominant frequencies alone are sufficient for accurate classification. An example of the results of these metrics at different threshold values are given in Fig. 3, where the size of each point reflects the number of classes in the two groups. Larger points indicate a higher number of classes, which suggests that the model has a stronger tendency to shortcut learning.
@@
Figure 3: ResNet18 tested on original test images and DFM-filtered images of ImageNet-1k. Blue line shows the average TPR of classes subject to shortcuts on the original test images, and the orange line shows the results of non-shortcut classes, at different threshold $t$. The green and red lines correspond to results tested on DFM-filtered images. Lower $t$ indicates weak shortcuts and higher $t$ signifies stronger ones. The size of each point reflects the number of classes. The larger the size, the more classes included.
@@ 3.3 Impact of shortcut on generalization
To assess how frequency shortcuts impact the generalization and robustness performance of models trained on ImageNet-1k, we evaluate their performance on several datasets. We perform ID tests on ImageNet-1k (IN-1k) [6] and ImageNet-v2 (IN-v2) [33] test sets. They are considered in-distribution as they are collected using the same protocol as the training set. Furthermore, we use ImageNet-C (IN-C) [15] to benchmark the robustness of models against appearance image corruptions, e.g. noise, blur and weather changes. To evaluate the generalizability of models to images with different renditions, we use ImageNet-Renditions (IN-R) [16] and ImageNet-Sketch (IN-S) [38]. IN-R contains renditions like cartoons, paintings, art and toys of 200 classes from IN-1k. IN-S has the same number of classes as IN-1k, containing images of hand-drawn sketches. These datasets contain visual renditions, which serve to evaluate how the reliance on texture cues (corresponding to most frequency shortcuts) impacts on generalization. To measure the adversarial robustness of models, we apply the fast gradient sign method (FGSM) attacks [10] to the validation set of IN-1k, with $L_{\infty}=4/255$. We calculate the average TPR of shortcut and non-shortcut classes on OOD data, that we compare to the results on ID data.
@@ 4.1 Configuration of HFSS
Figure 7: Average TPR of shortcut and non-shortcut classes given different thresholds on datasets with 1000 classes. Models generally perform better on images of shortcut classes than non-shortcut classes. This does not hold for results on IN-S, which lacks preserved texture information.
Figure 8: Average TPR of shortcut and non-shortcut classes at different thresholds on datasets with 200 classes. Models perform better on shortcut classes than non-shortcut classes in IN-200. IN-R lacks preserved texture information and thus the models have worse performance on shortcut classes than non-shortcut classes.
@@ 4.2 Shortcut identification on IN-1k
Different from the results on IN-v2 and IN-C, frequency shortcuts impair generalization of models to texture and rendition changes. Model performance on shortcut and non-shortcut classes on IN-S is close to each other. In Fig. 8, we observe that shortcut classes have worse prediction results than non-shortcut classes when tested on IN-R which contains images with rendition changes. This is attributable to the fact that shortcut information are not available in the OOD tests, as the rendition and sketch test sets preserve less or very different texture information than IN-C and IN-v2 (see green lines of IN-S and IN-1k , IN-R and IN-200 in Fig. 7 and Fig. 8). Frequency shortcut learning is an explicit cause of the texture-bias of ImageNet models, and the resulting impaired generalizability to renditions is in line with [8].
@@
Figure 9: An image of dugong contains frequency shortcuts which a model relies on for high-confidence classification. Its presence in OOD test sets like IN-C, and in the image under FGSM attacks yields correct predictions. But the reliance on shortcuts impairs model generalization performance to renditions changes.
@@ 4.3 Comparison with the existing approach
Frequency shortcuts can impair the generalization performance of models or provide a false impression of good generalization when shortcuts exist in OOD tests [41]. We thus validate the impact of shortcuts searched by HFSS in generalization tests, by measuring performance results on IN-SCT, an OOD test set designed by [41]. We report the results in Tab. 3. The strong shortcuts for classes airliner, frog and container ship searched by HFSS are present in the OOD set, contributing to close-to or above average TPR of classes Military aircraft (Mil-aircraft), tree frog, and fishing vessel in the IN-SCT dataset. The shortcuts searched by [41] for class Siamese cat result in a much lower TPR for class tabby cat in IN-SCT, though shortcuts for this class still exist in the OOD data (as observed from the results of HFSS). This indicates that the method proposed in [41] has limitations in estimating the impact of shortcuts on OOD performance, and further demonstrates the effectiveness of HFSS in finding shortcuts.
@@
Table 3: $\mathrm{TPR}$ results on IN-SCT (first row) and corresponding DFM-filtered images (second and third rows). $\mathrm{TPR}$ above average TPR (0.367) is highlighted in bold.